# supplement.pdf

### v2

@@ 2.2 Test Procedure
where $\sigma_{E}^{2}=\mathrm{Var}[E_{M,i}]=\mathrm{Var}[E_{i}^{\top}\theta_{M}]$ and $\sigma_{Z}^{2}=\mathrm{Var}[Z_{i}]$. In practice, an estimator $\hat{\sigma}_{Z}^{2}$ of $\sigma_{Z}^{2}$ could be derived from the scaled Lasso method. Let $\sigma^{2}$ be the variance of the residual when regressing $Y_{i}$ on $A_{i}$. In light of $\sigma_{E}^{2}=\sigma^{2}-\sigma_{Z}^{2}$, we can estimate $\sigma_{E}^{2}$ by $\hat{\sigma}_{E}^{2}=\max\{\hat{\sigma}^{2}-\hat{\sigma}_{Z}^{2},0\}$, where $\hat{\sigma}^{2}$ is the sample version of $\sigma^{2}$. This allows us to obtain an estimated variance $\hat{V}_{0}=\frac{\hat{\sigma}_{E}^{2}}{n}\hat{\Sigma}_{A}^{-1}+\frac{\hat{\sigma}_{Z}^{2}}{n}\hat{U}^{\top}\hat{\Sigma}_{X}\hat{U}$. When the null of (3) holds, the scenario where $\theta_{M}=0$ and $\beta_{A}=0$ is of particular interest as it is commonly encountered in genome-wide analyses (Huang, 2019), which may lead to a super-efficiency issue. To illustrate this, consider the case of $q=1$ for simplicity. Since $\sigma_{E}^{2}=0$ when $\theta_{M}=0$, the standard deviation $\sqrt{V_{0}}$ is reduced to $\sqrt{V_{0}}=\frac{\sigma_{Z}}{\sqrt{n}}(\hat{u}_{1}^{\top}\hat{\Sigma}_{X}\hat{u}_{1})^{1/2}$. The magnitude of this value is in the order of $n^{-1/2}\|g_{1}\|_{2}$, where $\|g_{1}\|_{2}=o_{P}(1)$ in certain cases when $\beta_{A}=0$, according to Lemma LABEL:lemma:bound-G_21 of Supplementary Material. Consequently, the magnitude of $\sqrt{V_{0}}$ converges to zero at the rate $n^{-1/2}$ when $\beta_{A}\neq 0$ but potentially at a rate faster than $n^{-1/2}$ when $\beta_{A}=0$, making it difficult to construct a valid test. To address this issue, we introduce a ridge to $V_{0}$ and consider the enlarged covariance matrix
@@
where $I_{q}$ is the $q\times q$ identity matrix, and $\tau>0$ is a positive constant; a similar strategy was adopted in Guo et al. (2021b) with recommended $\tau=0.5$ or $\tau=1$. Based on the debiased estimator $\hat{\gamma}$ from (12) and the corresponding estimated variance $\hat{V}$ from (14), we propose the test statistics $\|T\|_{\infty}$ with
@@
2.3 Incorporating Potential Confounders
To extend the proposed test to incorporate potential confounders, consider the following extended model,
$M$ $=A\beta_{A}^{\top}+C\beta_{C}^{\top}+E,$
$Y$ $=A\theta_{A}+C\theta_{C}+M\theta_{M}+Z,$
where $C\in\mathbb{R}^{n\times r}$ represents a matrix of potential confounders, and $\beta_{C}$ and $\theta_{C}$ are the corresponding coefficients. Taking $X=(A,C,M)\in\mathbb{R}^{n\times(q+r+p)}$ and $\theta=(\theta_{A}^{\top},\theta_{C}^{\top},\theta_{M}^{\top})^{\top}\in\mathbb{R}^{q+r+p}$, we can still rewrite the second equation by
$Y=X\theta+Z,$
in analogy to (6). Letting $\beta=(\beta_{A},\beta_{C})\in\mathbb{R}^{p\times(q+r)}$, we observe that the mediation effect $\gamma=\beta_{A}^{\top}\theta_{M}$ corresponds to the first $q$ components of $\beta^{\top}\theta_{M}$. With $H=(A,C)\in\mathbb{R}^{n\times(q+r)}$, the ordinary least-squares estimator $\hat{\beta}=((H^{\top}H)^{-1}H^{\top}M)^{\top}$ of $\beta$ and an initial estimator $\hat{\theta}=(\hat{\theta}_{A}^{\top},\hat{\theta}_{C}^{\top},\hat{\theta}_{M}^{\top})^{\top}$ of $\theta$ form a pilot estimator $\tilde{\gamma}=\hat{\beta}_{A}^{\top}\hat{\theta}_{M}$ for $\gamma$, where $\hat{\beta}_{A}$ represents the first $q$ columns of $\hat{\beta}$. Modifying the bias correction procedure outlined in Section 2.1 in a straightforward way, we can obtain a debiased estimator $\hat{\gamma}$ for $\gamma$ when potential confounders are present. The debiased estimator is given by
$g_{j}=(\textbf{0}_{q+r}^{\top},\tilde{g}_{j}^{\top})\in\mathbb{R}^{q+r+p},$
@@ 3.1 Assumptions
To state our assumptions, we first introduce the concepts of subGaussianity and norm-subGaussianity. A real-valued random variable $S$ is subGaussian with a parameter $\sigma>0$ if $\mathbb{P}(|S-\mathbb{E}S|\geq t)\leq 2e^{-t^{2}/(2\sigma^{2})}$. When $S\in\mathbb{R}^{p}$ is a random vector, it is subGaussian with a parameter $\sigma>0$ if $v^{\top}S$ is subGaussian with the parameter $\sigma$ for all $v\in\mathbb{R}^{p}$ such that $\|v\|_{2}=1$. A random vector $S\in\mathbb{R}^{p}$ is norm-subGaussian with a parameter $\sigma>0$ (Jin et al., 2019), if $\mathbb{P}\left(\|S-\mathbb{E}S\|_{2}\geq t\right)\leq 2e^{-t^{2}/(2\sigma^{2})}$ for all $t\in\mathbb{R}$. Norm-subGaussianity generalizes the usual subGaussianity as both subGaussian and bounded random vectors are norm-subGaussian (Lemma 1, Jin et al., 2019). Recall $X=(A,M)$, $\theta=(\theta_{A}^{\top},\theta_{M}^{\top})^{\top}$, $Z=(Z_{1},\ldots,Z_{n})^{\top}$, $E=(E_{1},\ldots,E_{n})^{\top}$, and $E_{M}=(E_{M,1},\ldots,E_{M,n})^{\top}=E\theta_{M}$ in (4) and (6). To study the theoretical properties of the proposed test, we require the following assumptions.
@@
    $\max_{1\leq i\leq n}\max\left\{\mathbb{E}[E_{M,i}^{2+\nu}|X_{i}],\mathbb{E}[Z_{i}^{2+\nu}|X_{i}]\right\}\leq M_{0}$ for some constants $\nu,M_{0}>0$.
Assumptions similar to (A1) are commonly adopted in the high-dimensional statistics and high-dimensional mediation analysis (Bühlmann and Van De Geer, 2011; Guo et al., 2022b). Assumption (A1) implies $\mathbb{E}[E_{M,i}|A_{i}]=0$ and $\mathbb{E}[E_{M,i}^{2}|A_{i}]=\sigma_{E}^{2}$ for some positive constant $0\leq\sigma_{E}^{2}<\infty$. We remark that no independence between $X_{i}$, $E_{i}$ and $Z_{i}$ is required. In contrast, Zhou et al. (2020) assumed independence between $E_{i}$ and $A_{i}$ and independence between $Z_{i}$ and $X_{i}$. In addition to these independence assumptions, Guo et al. (2022b) further assumed that $E_{i}$ and $Z_{i}$ are independent. In fact, our conditional moment assumptions on $E_{i}$ and $Z_{i}$ in Assumption (A1), satisfied under the independence assumptions of Zhou et al. (2020) and Guo et al. (2022b), can accommodate more general dependence structures among $X_{i}$, $E_{i}$ and $M_{i}$. Moreover, Assumption (A1) guarantees the existence of the ordinary least-squares estimator $\hat{\beta}_{A}$ with probability tending to one, even when $q$ grows with $n$; see the proof of Lemma LABEL:lemma:bound-G_21 in Supplementary Material for details. Assumption (A2) poses further mild moment conditions on $E_{M,i}$ and $Z_{i}$. We also require the following general assumptions on the initial estimators.
@@
The convergence of $\hat{\sigma}_{Z}^{2}$ in Assumption (B2) holds when $\sigma_{Z}^{2}$ is estimated by the scaled Lasso (Sun and Zhang, 2012). Assumption (B1) can be satisfied by Lasso-type estimators under the compatibility condition and some sparsity structures. To introduce the compatibility condition, given a set $\mathcal{S}\subset\{1,\ldots,q+p\}$, for any positive number $\eta\geq 1$, define the set
@@
where $|\mathcal{S}|$ is the cardinality of $\mathcal{S}$. We say a design matrix $X$ satisfies the compatibility condition over $\mathcal{S}$ with a parameter $\eta$ if $\phi_{0}(\hat{\Sigma},\mathcal{S},\eta)>0$, where $\hat{\Sigma}=n^{-1}X^{\top}X$ is the sample covariance matrix of $X$. Similarly, we say that a (population) covariance matrix $\Sigma$ satisfies the compatibility condition over $\mathcal{S}$ with a parameter $\eta$ if $\phi_{0}(\Sigma,\mathcal{S},\eta)>0$. It is known that the compatibility condition is implied by the restricted eigenvalue condition, which is another well-known condition used to establish the consistency of Lasso-type estimators (van de Geer and Bühlmann, 2009). For the sparsity structures, we consider the capped-$\ell_{1}$ sparse structure on the regression coefficient $\theta$,
@@
with $\lambda_{0}=\sqrt{2\log(q+p)/n}$, which has been extensively examined (Sun and Zhang, 2013; Zhang and Zhang, 2014; Cai et al., 2021). As highlighted in Zhang and Zhang (2014), the capped-$\ell_{1}$ condition in (17) holds when $\theta$ is $\ell_{0}$ sparse with $\|\theta\|_{0}\leq s$ or $\ell_{r}$ sparse with $\|\theta\|_{r}^{r}/(\sigma_{Z}\lambda_{0})^{r}\leq s$ for $0<r\leq 1$. The following proposition shows that Lasso-type estimators satisfy Assumption (B1).
@@
    The scaled Lasso estimators $(\hat{\theta},\hat{\sigma}_{Z})$ (Sun and Zhang, 2012) satisfy Assumptions (B1) and (B2).
@@ 3.2 Validity and Consistency
[…]
@@
that encodes a capped-$\ell_{1}$ condition only on the vector $\theta$, where $M_{0}>0$ and $C_{0}\geq c_{0}>0$ are positive constants defined in Assumptions (A1) and (A2), and $\lambda_{0}=\sqrt{2\log(q+p)/n}$. Recall $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution, and let $\mathbb{P}_{\xi}$ be the probability measure induced by the parameter $\xi$. Define $F(\alpha,x,q)=q\left\{\Phi\left(x+\Phi^{-1}\left(1-{\alpha}/{(2q)}\right)\right)-\Phi\left(x-\Phi^{-1}\left(1-{\alpha}/{(2q)}\right)\right)\right\}$ for any $x\in\mathbb{R}$, as well as $\delta_{\max}=\max_{1\leq k\leq q}|\delta_{k}|$ and $\beta_{A,\max}=\max_{1\leq k\leq q}\|\beta_{A,k}\|_{2}$.
@@
  • (Power). The power of the test $\phi_{\alpha}$ under the local alternative is asymptotically lower bounded by $1-F(\alpha,\Delta_{n},q)$, where $\Delta_{n}={C\delta_{\max}}/{({\sigma\sqrt{\log(pn)/n}}+\beta_{A,\max}+C)}$ for some constant $C>0$. That is,
@@
In contrast with Zhou et al. (2020); Guo et al. (2022b) that require additional assumptions on $\beta_{A}$, Theorem 2 shows that the proposed test remains valid for arbitrary $\beta_{A}$, as the parameter space $\Xi$ encodes no assumptions on $\beta_{A}$.
@@ 4 Simulation Study
In this section, we conduct numerical simulations to assess the performance of the proposed test $\phi_{\alpha}$ (denoted as Proposed) described in Section 2.2. We employ the Lasso estimator as the initial estimator for $\theta$ and select the related tuning parameter by the quantile-based penalty procedure in the R package scalreg (Sun, 2019). The tuning parameter $\lambda$ in (9) and (10) is chosen by the searching procedure in the R package SIHR (Rakshit et al., 2022). We set $\tau=1$ in (14). Throughout the simulations, we consider the model specified by (1) and (2) as our data generating model, and adopt simulation settings similar to those of Zhou et al. (2020) and Guo et al. (2022b).
@@ 4.1 Size
The null hypothesis of (3) is composed of the following four cases: 1) $\beta_{A}=0$ and $\theta_{M}=0$, 2) $\beta_{A}\neq 0$ and $\theta_{M}=0$, 3) $\beta_{A}=0$ and $\theta_{M}\neq 0$ and 4) $\beta_{A}^{\top}\theta_{M}=0$ but $\beta_{A}\neq 0$ and $\theta_{M}\neq 0$. In this section, we numerically assess the validity of the proposed test procedure and the alternative existing methods for each of the these cases. We consider $n=50$ and $n=300$ which correspond to small and large sample sizes. Set $p=n$ and consider two cases of $q$, namely, $q=1$ and $q=3$. When the true coefficient vector $\theta_{M}\in\mathbb{R}^{p}$ is nonzero, three types of sparsity similar to those of Cai et al. (2021) are considered with a fixed sparsity parameter $s=5$.
  • Hard sparsity: $\theta_{M,k}=0.2k\mathbb{I}_{\left\{1\leq k\leq s\right\}}$, for $k=1,\ldots,p$;
When the true coefficient matrix $\beta_{A}\in\mathbb{R}^{p\times q}$ is nonzero, each of its elements is given by
@@
for $j=1,\ldots,q$, where each $\kappa_{j}(\cdot)$ is a random permutation of $\{1,\ldots,s\}$. For the coefficient $\theta_{A}$, we set $\theta_{A}=c_{0}\textbf{1}_{q}$ with $c_{0}=0.5$ for $q=1$ and $c_{0}=0.3$ for $q=3$. For the exposure $A_{i}=(A_{ij})_{j=1}^{q}$, the coordinates of $A_{i}$ are independently drawn from a distribution, for which we consider two cases: The Bernoulli distribution $\text{Bern}(0.5)$ and the Gaussian distribution $N(0,0.5^{2})$. Set $Z_{i}\sim N(0,1)$ and $E_{i}\sim N_{q}(0,\Sigma_{0})$ for $\Sigma_{0}=\Sigma_{AR}=(0.5^{|i-j|})_{1\leq i,j\leq p}$. For each experimental setting, we independently generate $R=500$ replications. For each replication, the empirical size is computed as the proportion of rejections among the $R$ replications when the composite null hypothesis of (3) is true. Table 2 presents the numerical results comparing the proposed procedure with the methods of Zhou et al. (2020) (referred to as Zhou2020) and Guo et al. (2022b) (referred to as Guo2022) that also address the overall mediation effect $\gamma$ and the test problem (3), for the case when the exposures follow the Bernoulli distribution $\text{Bern}(0.5)$; the results for Gaussian exposures are provided in Supplementary Material. The tuning parameters of Zhou2020 and Guo2022 were respectively selected via the R package developed by Zhou et al. (2020) and the method described in Guo et al. (2022b). The results show that the proposed test procedure remains valid in all cases of the composite null hypothesis of (3) and in all settings of $\beta_{A}$. In contrast, Zhou2020 appears to be invalid when $\theta_{M}=0$, particularly when both $\beta_{M}=0$ and $\beta_{A}=0$, a case that is not covered by the theory of Zhou et al. (2020). The method Guo2022 seems invalid when the sample size is small, for example, when $n=50$. Interestingly, although the theoretical analysis in Guo et al. (2022b) requires a uniform signal strength condition for $\theta_{M}$, that is, $\theta_{M}$ should be sufficiently large, the method Guo2022 appears to be conservative when $\Sigma_{0}=\Sigma_{AR}$, $\theta_{M}=0$ and $n$ is relatively large. This phenomenon might be attributed to the relatively aggressively selected tuning parameter that tends to yield a zero estimator of $\theta_{M}$ in the case of $\theta_{M}=0$. Since the irrepresentable condition, on which the theory of Guo et al. (2022b) essentially relies, is usually not satisfied when covariates in a high-dimensional linear model (such as mediators in the LSEM) are highly correlated (Zhao and Yu, 2006; Ravikumar et al., 2010), it is of interest to also investigate the behavior of the tests when the irrepresentable condition might fail. For this purpose, in the case of $\theta_{M}=0$, we consider an additional scenario $\Sigma_{0}=\Sigma_{CS}$ for a compound symmetric covariance matrix $\Sigma_{CS}$, where the diagonal elements of $\Sigma_{CS}$ are $8$ and the other elements are $6.4$. This choice of $\Sigma_{CS}$ corresponds to a more challenging setting where the mediators exhibit strong correlations. The results, also presented in Table 2, suggest that the method Guo2022 appears to become invalid even when $n=300$, while our test remains valid. We would like to remark that the results presented above do not suggest the failure of the existing methods. Rather, they underscore that these methods require stronger conditions to operate effectively. When these conditions are met, these methods often exhibit commendable power.

### v3

@@ 2.2 Test Procedure
where $\sigma_{E}^{2}=\mathrm{Var}[E_{M,i}]=\mathrm{Var}[E_{i}^{\top}\theta_{M}]$ and $\sigma_{Z}^{2}=\mathrm{Var}[Z_{i}]$. In practice, an estimator $\hat{\sigma}_{Z}^{2}$ of $\sigma_{Z}^{2}$ could be derived from the scaled Lasso method. Let $\sigma^{2}$ be the variance of the residual when regressing $Y_{i}$ on $A_{i}$. In light of $\sigma_{E}^{2}=\sigma^{2}-\sigma_{Z}^{2}$, we can estimate $\sigma_{E}^{2}$ by $\hat{\sigma}_{E}^{2}=\max\{\hat{\sigma}^{2}-\hat{\sigma}_{Z}^{2},0\}$, where $\hat{\sigma}^{2}$ is the sample version of $\sigma^{2}$. This allows us to obtain an estimated variance $\hat{V}_{0}=\frac{\hat{\sigma}_{E}^{2}}{n}\hat{\Sigma}_{A}^{-1}+\frac{\hat{\sigma}_{Z}^{2}}{n}\hat{U}^{\top}\hat{\Sigma}_{X}\hat{U}$.
When the null of (3) holds, the scenario where $\theta_{M}=0$ and $\beta_{A}=0$ is of particular interest as it is commonly encountered in genome-wide analyses [10], which may lead to a super-efficiency issue. To illustrate this, consider the case of $q=1$ for simplicity. Since $\sigma_{E}^{2}=0$ when $\theta_{M}=0$, the standard deviation $\sqrt{V_{0}}$ is reduced to $\sqrt{V_{0}}=\frac{\sigma_{Z}}{\sqrt{n}}(\hat{u}_{1}^{\top}\hat{\Sigma}_{X}\hat{u}_{1})^{1/2}$. The magnitude of this value is in the order of $n^{-1/2}\|g_{1}\|_{2}$, where $\|g_{1}\|_{2}=o_{P}(1)$ in certain cases when $\beta_{A}=0$, according to Lemma 1 in Appendix D. Consequently, the magnitude of $\sqrt{V_{0}}$ converges to zero at the rate $n^{-1/2}$ when $\beta_{A}\neq 0$ but potentially at a rate faster than $n^{-1/2}$ when $\beta_{A}=0$, making it difficult to find an accuracy estimator of $\sqrt{V_{0}}$ for constructing a valid test. To address this issue, we introduce a ridge to $V_{0}$ and consider the enlarged covariance matrix
@@
where $I_{q}$ is the $q\times q$ identity matrix, and $\tau>0$ is a positive constant; a similar strategy was adopted in [25] with recommended $\tau=0.5$ or $\tau=1$.
Based on the debiased estimator $\hat{\gamma}$ from (12) and the corresponding estimated variance $\hat{V}$ from (14), we propose the test statistics $\|T\|_{\infty}$ with
@@
Remark 1.
An extension for incorporating additional covariates is provided in Appendix B.
@@ 3.1 Assumptions
To state our assumptions, we first introduce the concepts of subGaussianity and norm-subGaussianity. A real-valued random variable $S$ is subGaussian with a parameter $\sigma>0$ if $\mathbb{P}(|S-\mathbb{E}S|\geq t)\leq 2e^{-t^{2}/(2\sigma^{2})}$. When $S\in\mathbb{R}^{p}$ is a random vector, it is subGaussian with a parameter $\sigma>0$ if $v^{\top}S$ is subGaussian with the parameter $\sigma$ for all $v\in\mathbb{R}^{p}$ such that $\|v\|_{2}=1$. A random vector $S\in\mathbb{R}^{p}$ is norm-subGaussian with a parameter $\sigma>0$ [26], if $\mathbb{P}\left(\|S-\mathbb{E}S\|_{2}\geq t\right)\leq 2e^{-t^{2}/(2\sigma^{2})}$ for all $t\in\mathbb{R}$. Norm-subGaussianity generalizes the usual subGaussianity as both subGaussian and bounded random vectors are norm-subGaussian [Lemma 1, 26]. Recall $X=(A,M)$, $\theta=(\theta_{A}^{\top},\theta_{M}^{\top})^{\top}$, $Z=(Z_{1},\ldots,Z_{n})^{\top}$, $E=(E_{1},\ldots,E_{n})^{\top}$, and $E_{M}=(E_{M,1},\ldots,E_{M,n})^{\top}=E\theta_{M}$ in (4) and (6). To study the theoretical properties of the proposed test, we require the following assumptions.
@@
    $\max_{1\leq i\leq n}\max\left\{\mathbb{E}[E_{M,i}^{2+\nu}|X_{i}],\mathbb{E}[Z_{i}^{2+\nu}|X_{i}]\right\}\leq M_{0}$ for constants $\nu,M_{0}>0$.
Assumptions similar to (A1) are commonly adopted in the high-dimensional statistics and high-dimensional mediation analysis [27, 17]. Assumption (A1) implies $\mathbb{E}[E_{M,i}|A_{i}]=0$ and $\mathbb{E}[E_{M,i}^{2}|A_{i}]=\sigma_{E}^{2}$ for some positive constant $0\leq\sigma_{E}^{2}<\infty$. We remark that no independence between $X_{i}$, $E_{i}$ and $Z_{i}$ is required. In contrast, [16] assumed independence between $E_{i}$ and $A_{i}$ and independence between $Z_{i}$ and $X_{i}$. In addition to these independence assumptions, [17] further assumed that $E_{i}$ and $Z_{i}$ are independent. In fact, our conditional moment assumptions on $E_{i}$ and $Z_{i}$ in Assumption (A1), satisfied under the independence assumptions of [16] and [17], can accommodate more general dependence structures among $X_{i}$, $E_{i}$ and $M_{i}$. Moreover, Assumption (A1) guarantees the existence of the ordinary least-squares estimator $\hat{\beta}_{A}$ with probability tending to one, even when $q$ grows with $n$; see the proof of Lemma 1 in Appendix D for details. Assumption (A2) poses further mild moment conditions on $E_{M,i}$ and $Z_{i}$.
We also assume the following conditions on the initial estimators.
@@
The convergence of $\hat{\sigma}_{Z}^{2}$ in Assumption (B2) holds when $\sigma_{Z}^{2}$ is estimated by the scaled Lasso [28].
Assumption (B1) can be satisfied by Lasso-type estimators under the compatibility condition and certain sparsity structures, where $s$ serves as a sparse parameter. To introduce the compatibility condition, given a set $\mathcal{S}\subset\{1,\ldots,q+p\}$, for any positive number $\eta\geq 1$, define the set
@@
where $|\mathcal{S}|$ is the cardinality of $\mathcal{S}$. We say a design matrix $X$ satisfies the compatibility condition over $\mathcal{S}$ with a parameter $\eta$ if $\phi_{0}(\hat{\Sigma},\mathcal{S},\eta)>0$, where $\hat{\Sigma}=n^{-1}X^{\top}X$ is the sample covariance matrix of $X$. Similarly, we say that a (population) covariance matrix $\Sigma$ satisfies the compatibility condition over $\mathcal{S}$ with a parameter $\eta$ if $\phi_{0}(\Sigma,\mathcal{S},\eta)>0$. It is known that the compatibility condition is implied by the restricted eigenvalue condition, which is another well-known condition used to establish the consistency of Lasso-type estimators [29].
For the sparsity structures, in this paper, we consider the capped-$\ell_{1}$ sparse structure on the regression coefficient $\theta$,
@@
with $\lambda_{0}=\sqrt{2\log(q+p)/n}$, which has been extensively examined [30, 18, 22]. As highlighted in [18], the capped-$\ell_{1}$ condition in (17) holds when $\theta$ is $\ell_{0}$ sparse with $\|\theta\|_{0}\leq s$ or $\ell_{r}$ sparse with $\|\theta\|_{r}^{r}/(\sigma_{Z}\lambda_{0})^{r}\leq s$ for $0<r\leq 1$. The following proposition shows that Lasso-type estimators satisfy Assumption (B1).
@@
    The scaled Lasso estimators $(\hat{\theta},\hat{\sigma}_{Z})$ [28] satisfy Assumptions (B1) and (B2).
@@ 3.2 Validity and Consistency
$\hat{\gamma}-\gamma=W+B,\quad\text{with}\quad W_{j}/\sqrt{V_{jj}}\stackrel{{\scriptstyle d}}{{\to}}N(0,1)~{}\text{for each}~{}j=1,\ldots,q,$
@@
that encodes a capped-$\ell_{1}$ condition only on the vector $\theta$, where $M_{0}>0$ and $C_{0}\geq c_{0}>0$ are positive constants defined in Assumptions (A1) and (A2), and $\lambda_{0}=\sqrt{2\log(q+p)/n}$. Recall $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution, and let $\mathbb{P}_{\xi}$ be the probability measure induced by the parameter $\xi$. Define
$F(\alpha,x,q)=q\left\{\Phi\left(x+\Phi^{-1}\left(1-{\alpha}/{(2q)}\right)\right)-\Phi\left(x-\Phi^{-1}\left(1-{\alpha}/{(2q)}\right)\right)\right\}$
for any $x\in\mathbb{R}$, as well as $\delta_{\max}=\max_{1\leq k\leq q}|\delta_{k}|$ and $\beta_{A,\max}=\max_{1\leq k\leq q}\|\beta_{A,k}\|_{2}$.
@@
  • (Power). The power of the test $\phi_{\alpha}$ under the local alternative is asymptotically lower bounded by $1-F(\alpha,\Delta_{n},q)$, where
    $\Delta_{n}={C\delta_{\max}}/{({\sigma\sqrt{\log(pn)/n}}+\beta_{A,\max}+C)}$
    for some constant $C>0$. That is,
@@
In contrast with [16, 17] that require additional assumptions on $\beta_{A}$, Theorem 3 shows that the proposed test remains valid for arbitrary $\beta_{A}$, as the parameter space $\Xi$ encodes no assumptions on $\beta_{A}$. While $q$ (the number of exposures) is allowed to grow with $n$ in Theorem 3, in practice the number of exposures or treatments is often low-dimensional. Note that for any fixed (or sufficiently slowly diverging) $q\geq 1$ and $\alpha>0$, the function $F(\alpha,x,q)$ is monotone decreasing with respect to $x$. That is, a larger value of $\Delta_{n}$ results in a higher power for the proposed test $\phi_{\alpha}$. Furthermore, if mediation effect is of constant order, i.e., $\|\beta_{A}^{\top}\theta_{M}\|_{\infty}=O(1)$, then $\delta_{\max}\asymp\sqrt{n}$. This, together with constant-order exposure-mediator coefficients ($\beta_{A,\max}=O(1)$), further implies that $\Delta_{n}\to\infty$ as $n\to\infty$. Consequently, the proposed test $\phi_{\alpha}$ achieves asymptotic power one.
@@ 4 Simulation Study
In this section, we conduct numerical simulations to assess the performance of the proposed test $\phi_{\alpha}$ (denoted as Bonf-$1$) described in Section 2.2. We employ the scaled Lasso estimator as the initial estimator for $\theta$ and select the related tuning parameter by the quantile-based penalty procedure in the R package scalreg [31]. The tuning parameter $\lambda$ in (9) and (10) is chosen by the searching procedure in the R package SIHR [32]. We set $\tau=1$ in (14). Throughout the simulations, we consider the model specified by (1) and (2) as our data generating model, and adopt simulation settings similar to those of [16] and [17].
@@ 4.1 Size
The null hypothesis of (3) is composed of the following four cases: 1) $\beta_{A}=0$ and $\theta_{M}=0$, 2) $\beta_{A}\neq 0$ and $\theta_{M}=0$, 3) $\beta_{A}=0$ and $\theta_{M}\neq 0$ and 4) $\beta_{A}^{\top}\theta_{M}=0$ but $\beta_{A}\neq 0$ and $\theta_{M}\neq 0$. In this section, we numerically assess the validity of the proposed test procedure and the alternative existing methods for each of the these cases.
We consider $n=50$ and $n=300$ which correspond to small and large sample sizes, respectively. We set $p=n$ and consider two cases for $q$: $q=1$ and $q=3$. When the true coefficient vector $\theta_{M}\in\mathbb{R}^{p}$ is nonzero, we investigate a hard sparsity case with a fixed sparsity parameter $s=5$: $\theta_{M,k}=0.2k\mathbb{I}_{\left\{1\leq k\leq s\right\}}$ for $k=1,\ldots,p$. We also explore other types of sparsity in Appendix E. When the true coefficient matrix $\beta_{A}\in\mathbb{R}^{p\times q}$ is nonzero, we consider two scenarios: a sparse setting and a dense setting. In the sparse setting, each element of $\beta_{A}$ is given by
@@
for $s=5$, $k=1,\ldots,p$ and $j=1,\ldots,q$, where each $\kappa_{j}(\cdot)$ is a random permutation of $\{1,\ldots,s\}$. In the dense setting, each element of $\beta_{A}$ is given by
$\beta_{A,jk}=\begin{cases}0.2&\text{ if }s+1\leq k\leq p/2\\ 0&\text{ otherwise,}\end{cases}$
for $s=5$, $k=1,\ldots,p$ and $j=1,\ldots,q$. For the coefficient $\theta_{A}$, we set $\theta_{A}=c_{0}\textbf{1}_{q}$ with $c_{0}=0.5$ for $q=1$ and $c_{0}=0.3$ for $q=3$.
For the exposure $A_{i}=(A_{ij})_{j=1}^{q}$, the coordinates of $A_{i}$ are independently drawn from a distribution, for which we consider two cases: The Bernoulli distribution $\text{Bern}(0.5)$ and the Gaussian distribution $N(0,0.5^{2})$. For error terms, we set $Z_{i}\sim N(0,1)$ and $E_{i}\sim N_{q}(0,\Sigma_{0})$ with two scenarios for the covariance structure. In the first scenario, we use $\Sigma_{0}=\Sigma_{AR}=(0.5^{|i-j|})_{1\leq i,j\leq p}$, representing a covariance matrix with an AR(1) structure. In the second scenario, we consider $\Sigma_{0}=\Sigma_{CS}$ for a compound symmetric covariance matrix $\Sigma_{CS}$, where the diagonal elements of $\Sigma_{CS}$ are $8$ and the other elements are $6.4$. This choice of $\Sigma_{CS}$ is a more challenging setting where the mediators exhibit strong correlations, corresponding to a scenario where the irrepresentable condition may fail [19]. For each experimental setting, we independently generate $R=500$ replications. For each replication, the empirical size is computed as the proportion of rejections among the $R$ replications when the composite null hypothesis of (3) is true.
For comprehensive comparisons, we consider several competitors. First, to assess the effect of $\tau$ in the proposed covariance matrix (14), we introduce a test, denoted Bonf-00, which is similar to Bonf-$1$ but with $\tau=0$. Next, recall that our proposed test (15) is based on the Bonferroni correction. Another possible approach is to perform a $\chi^{2}$ test using the de-biased estimator $\hat{\gamma}$ from (12) and the corresponding estimated variance $\hat{V}$ from (14). We thus consider the $\chi^{2}$ tests with $\tau=0$ and $\tau=1$ in (14), referred to as $\chi^{2}$-00 and $\chi^{2}$-$1$, respectively. Note that when $q=1$, the proposed Bonferroni-based test (15) and this $\chi^{2}$ test are equivalent. Additionally, we include the methods of [16] (denoted Zhou-00) and [17] (denoted Guo-00), which also address the overall mediation effect $\gamma$ and the test problem (3). The tuning parameters for Zhou-00 and Guo-00 were selected using the R package developed by [16] and the method described in [17], respectively. To examine whether the existing methods can benefit from the enlarged covariance strategy, as in $\hat{V}$ from (14), we also investigate tests in [16] and [17] based on the covariance matrices with an additional enlarged factor $1/n\cdot I_{q}$ as in (14), referred to as Zhou-$1$ and Guo-$1$, respectively.
@@
Table 2 presents the numerical results for the case when the exposures follow the Bernoulli distribution $\text{Bern}(0.5)$; the results for Gaussian exposures are provided in Appendix E. The results show that the proposed test, Bonf-$1$, remains valid across all cases of the composite null hypothesis in (3) and for all settings of $\beta_{A}$. In contrast, Zhou-00 and Zhou-$1$ are invalid when $\theta_{M}=0$, while Guo-00 and Guo-$1$ tend to fail when the sample size is small ($n=50$) or mediators are highly correlated ($\Sigma_{0}=\Sigma_{\text{CS}}$). Interestingly, although the theoretical analysis in [17] requires a uniform signal strength condition for $\theta_{M}$ (i.e., $\theta_{M}$ should be sufficiently large), the method Guo-00 and Guo-$1$ appear to be conservative when $\Sigma_{0}=\Sigma_{AR}$, $\theta_{M}=0$ and $n$ is relatively large. This phenomenon might be attributed to the relatively aggressively selected tuning parameter that tends to yield a zero estimator of $\theta_{M}$ when $\theta_{M}=0$. We would like to remark that the results presented above do not suggest the failure of the existing methods. Rather, they underscore that these methods require stronger conditions to operate effectively. When these conditions are met, these methods often exhibit commendable power.
In addition, the results in Table 2 provide further insights. First, the proposed test benefits primarily from two key components: the VePD de-biasing technique and the regularization $\tau/n\cdot I_{q}$ of the variance. Comparing the results of Bonf-$1$ with those of Bonf-00, the impact of $\tau$ appears to be relatively modest. Furthermore, Zhou-$1$ (Guo-$1$) shows only a slight adjustment over Zhou-00 (Guo-00). These findings suggest that the proposed test’s effectiveness largely stems from the VePD technique, which differs fundamentally from the methods of [16] and [17]. Second, the performance of the proposed Bonferroni-based test (Bonf-$1$) is comparable to the $\chi^{2}$ test ($\chi^{2}$-$1$) in most simulation scenarios. This indicates that the proposed Bonferroni-based test is not overly conservative, even when compared to a method that accounts for the correlation structure of the de-biased estimator.